OpenStudy (anonymous):

Factor completley. 6x^2+20x+16.

OpenStudy (zpupster):

you can divide all terms by 2

OpenStudy (anonymous):

not GCF, factor the trinomial.

OpenStudy (nirmalnema):

simplify it... you get 3x^2+10x+8

OpenStudy (anonymous):

A(2x+1)(3x-5) B.(3x+1)(3x-6) C.(6x+1)(4x-1) D.(x-3)(6x-4)

OpenStudy (zpupster):

then factor ok just multiply what you have then using foil method

OpenStudy (nirmalnema):

3x^2+6x+4x+8 3x(x+2)+4(x+2)

OpenStudy (anonymous):

....nevermind, that doesnt really help, to factor trinomials is different from factoring a regular algebraic equation you have to take the middle term, and find two number that are equal substitute them......but that all i know how to do. Thank you for your help though.

OpenStudy (zpupster):

weird none of your las terms multiplied = 16 or 8 anyway 2(3x +4)(x + 2)

OpenStudy (zarkon):

none of the 'answers' you listed are the correct answer

OpenStudy (nirmalnema):

yup ......@Zarkon

OpenStudy (nirmalnema):

(3x+4)(x+2) the correct ans is this

OpenStudy (zpupster):

weird none of your las terms multiplied = 16 or 8 anyway 2(3x +4)(x + 2) no --this is the correct answer

OpenStudy (zarkon):

zpupster has the correct answer

OpenStudy (nirmalnema):

but both the ans are correct

OpenStudy (anonymous):

the correct answer according the the answer key is (3x+4)(2x+4)

OpenStudy (anonymous):

and sorry i posted the wrong choices.

OpenStudy (zarkon):

that is equivalent to zpupster's answer

OpenStudy (nirmalnema):

yup..

OpenStudy (zarkon):

\[2(3x +4)(x + 2)=(3x +4)2(x + 2)=(3x +4)(2\cdot x + 2\cdot2)\] \[(3x +4)(2x + 4)\]